The following chemical reaction takes place:
3 (NH4)2S(aq) + 2 AlCl3(aq) = Al2S3(s) + 6 NH3(g) + 6HCl(aq)
If a 400 mL solution of AlCl3 with a concentration of 1.10 mol/L is reacted with 3.56g (NH4)2S:
i) What is the concentration of HCl?
ii) How much Al2S3(s) is produced?
iii) If this reaction is carried at SATP conditions, what volume of NH3 is produced?
1
Expert's answer
2020-07-29T06:38:59-0400

First the limiting reactant should be determined.

n (AlCl3) = 1.10M x 0.4L = 0.44 mol

n (NH4)2S = 3.56g / 68.15g/mol = 0.0522 mol - evidently limiting reactant.


1) n (HCl) = 6/3 x n(NH4)2S = 0.1044 mol

[HCl] = 0.1044mol / 0.4L = 0.261 M


2) n (Al2S3) = 1/3 x n(NH4)2S = 0.0174 mol

m (Al2S3) = 0.0174mol x 150.16g/mol = 2.61 g


3) n (NH3) = 6/3 x n(NH4)2S = 0.1044 mol

V (NH3) = 0.1044mol x 22.4L/mol = 2.34 L